OpenStudy (anonymous):

Which expression is the simplified form of the square root of k to the seventeenth power.?

OpenStudy (anonymous):

\[\large \sqrt{k^{17}} = \sqrt{\color{blue}{\underline{k^{8} \times k^{8}}} \times k} = ??\]

OpenStudy (anonymous):

You can put out one \(k^{8}\) here out of the square root brackets.. After that you are left with only \(k\) in the square root brackets.. Can you do now ??

OpenStudy (anonymous):

*pull out..

OpenStudy (kainui):

\[(k^{1/2})=\sqrt{k}\]

OpenStudy (anonymous):

Thanks Guys!! :))

OpenStudy (swissgirl):

so what is the answer @luvbouncer11 ?

OpenStudy (anonymous):

\[k ^{8}\sqrt{k}\] ??

OpenStudy (swissgirl):

yupppppp :DDD

OpenStudy (anonymous):

thankkkk youuuuuu